OpenStudy (anonymous):

Need help, please Show that the set of symmetric polynomials in 3 variables is a subring of all polynomials in 3 variables.

OpenStudy (anonymous):

I need to arrange the form of symmetric polynomial in general. I can handle the proof. :)

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

My attempt: let S ={ symmetric polynomials in 3 variables} P={ polynomials in 3 variables} then \(S\subset P\) and we know that P is a ring, hence to prove S is a subring, we need prove: a) S is closed under addition and additive inverse b) S is closed under multiplication both a, b are not hard to do, just have problem with generalize the form of symmetric polynomials in 3 variables. :)

OpenStudy (anonymous):

Let \(f,g\in S\) then we have 6 forms of f, g they are all equal. They are f(x,y,z) = f(x, z, y)=f(y, x, z)= f(y, z, x)= f(z, x, y)= f(z, y, x) g(x,y,z) =g(x, z, y)=g(y, x, z)= g(y, z, x)= g(z, x, y)= g(z, y, x)

OpenStudy (anonymous):

Do I have to expand them ( if I do, how?) or just let them as they are and start proving?

OpenStudy (xapproachesinfinity):

damn rings heheh